OpenStudy (meilendurcer):

Dritown received 0.135 inches of rain in the spring of 2011. Write this decimal as a fraction in simplest form. A.7/50 B.1/8 C.13/100 D.27/200

OpenStudy (anonymous):

its D

OpenStudy (anonymous):

sorry its c or D

OpenStudy (anonymous):

it D check calculators

OpenStudy (anonymous):

27 divided by 200 +0.135

OpenStudy (anonymous):

sorry =0.135

OpenStudy (anonymous):

0.135/1000=27/200 0.135/5=27 1000/5=200 27/200

OpenStudy (anonymous):

correct u see